OpenStudy (anonymous):

A train approaching a crossing changes speed from 25 m/s to 10 m/s in 240 s. How can the train's acceleration be described?

OpenStudy (anonymous):

v = u + at is the formula you need. Here, v is the final velocity which in the question is given as 10 m/s u is the initial velocity which in the question is given as 25 m/s and t is time which is equal to 240 m/s. Now since none of the units need to be converted or changed so we fill in these values in the above mentioned equation. Thus we get : 10 = 25 +240a => (-15) = 240a => (-15)/240 = a => -1/16 = a Thus acceleration a = -1/16 m/s^2 Now since a comes out to be -ive hence it is retardation not acceleration.